# Differentiable Physics-based System Identification for Robotic Manipulation of Elastoplastic Materials

### v1

@@ 4.4 Generalisation
Figure 8: Visualising the real long-horizon manipulation motions and their simulations based on the recovered parameters. Every four rows display the real and the top-three best simulations of a motion-object pair.
@@
To recall, there are three unseen motions. The triple-poking motion moves the round end-effector to poke the object at three different locations along the $y$ axis. The flattening motion moves the cylinder end-effector to press and flatten the object towards the positive and negative $x$ directions. The poking-180-rotating motion moves the rectangle end-effector to press the object down and then rotate $180$ degrees about the $z$ axis. The triple-poking motion is less relevant to the friction coefficients, while the last two motions involve complex frictional contacts. Table 5 summarises the heightmap distances of the simulated results of the unseen motions (each with two object configurations). Figure 8 visualises the trajectories of the real and the top three best simulation trajectories. The animation videos of them are included in Supplementary materials. Figure 9 presents the resultant particles and their heightmaps of each motion and object configuration. The following observations can be drawn.
@@
First, in Figure 8, (as well as the animation video), it can be seen that, for all motion and object configuration pairs, the real and simulated trajectories are highly similar. Figure 9 also shows that, although discrepancies exist, the simulated particles and their heightmaps of the manipulation results are also highly similar across the loss functions and datasets. This visual similarity proves that DPSI can indeed generalise to unseen, longer, and more complex manipulation motions, exhibiting robustness against perception and data challenges. Moreover, we note that it is the single-datapoint cases that achieve two-thirds of the top-three lowest heightmap losses across the three unseen motions as shown in Table 5.
@@
Thirdly, the found solutions are not globally optimal. Despite the visual similarity, Table 5 shows that the best generalisation results are not always achieved by the cases with the best in-distribution validation performances, as the underlined and bolded texts rarely coincide. Qualitatively, from the heightmaps and trajectories, one can also recognise certain visual discrepancies between the real objects and simulations. This indicates a certain degree of discrepancy between the found parameters and the actual solutions. Therefore, it shows that the data (and the motions that produce them) used for optimisation in our experiments are insufficient. Although multiple causes could be possible, it is out of the scope of this work to design better data collection processes for finding the optimal parameters.
@@
Finally, the modelling limitations observed in the in-distribution visualisation still exist in the unseen motion visualisations (Figs. 8 and 9). The insufficient elastic returning at contact areas especially with sharp manipulator edges can be seen from all visualisations where the primary contact areas of the object rarely return to heights close to the real object. Floating displaced particles can be observed from the poking-180-rotating motion visualisations where many particles are carried away during rotations. The insufficient spreading of the influence of contacts can be recognised from the first object configuration of the flattening motion visualisations where the object is not tilted enough in the second half of the trajectory.
@@
In summary, the results and discussions above strongly support the argument that DPSI is indeed capable of predicting unseen, longer, and more complex deformable object manipulation dynamics by optimising with very few real-world noisy and incomplete point-cloud data collected by simple and short interacting motions.
@@
Table 4: Computational costs (time and GPU memory comsuption) of the TaiChi-based differentiable simulator used in the DPSI framework with a Nvidia GeForce RX 4090 GPU. The costs of simulations with three motions and particle densities are reported. CP.: compile-time costs. The first time a program is run takes extra time to compile TaiChi kernels. $100$ FB-iterations (min): the runtime of $100$ forward and backward computations in minutes.
Motion Object Num. datapoint 12 6 1 (rec.) 1 (round) 1 (cyl.)
PCD CD 6178.32 6089.17 6418.81 6620.95 6146.67
Config. PRT CD 6238.82 6256.54 6166.92 6167.83 5964.65
0 PCD EMD 6105.29 5997.06 6034.08 6042.02 6157.98
Triple- PRT EMD 6499.63 6201.23 6193.93 6478.28 6243.00
poking PCD CD 9860.21 10189.94 10288.10 10215.02 10085.14
Config. PRT CD 10247.12 10267.20 10323.14 10228.00 $9$.92
1 PCD EMD 10378.19 10539.97 10300.59 10345.66 9768.11
PRT EMD 10715.33 9735.01 10374.79 10320.68 9903.31
PCD CD 5241.38 5270.79 4777.82 5015.08 5186.37
Config. PRT CD 5241.58 5145.21 5306.65 5388.45 5100.66
0 PCD EMD 5465.86 4698.51 5678.32 5658.92 5405.66
Flattening PRT EMD 5667.67 5108.31 6190.99 5887.79 5208.29
PCD CD 10890.47 10452.40 9600.17 9688.61 11074.55
Config. PRT CD 11897.35 11450.01 10934.94 10692.00 12251.92
1 PCD EMD 10985.88 10353.19 12189.07 12421.18 10476.08
PRT EMD 10836.55 10977.26 12123.76 11286.20 10990.40
PCD CD 5161.34 5085.54 5302.23 7091.04 4979.64
Config. PRT CD 6900.43 6342.54 5024.69 5018.13 8079.13
0 PCD EMD 6321.36 5288.02 6467.52 7036.70 4981.49
Poking-180 PRT EMD 5447.25 5051.74 7483.40 5844.29 5188.73
rotating PCD CD 9436.61 10015.51 10034.32 9212.14 10396.27
Config. PRT CD 8577.18 8613.41 10461.23 10385.83 8268.34
1 PCD EMD 8585.16 9126.36 8745.00 8401.15 9929.82
PRT EMD 8800.82 9394.39 9102.43 8765.50 9583.31
Figure 9: Visualisations of the particles and their heightmaps after applying the unseen motions, simulated with the best set of physics parameters found at the level 2 contact complexity experiments. Darker colours of the heightmaps indicate greater heights of the object. For each motion, there are two object configurations. The results for each trajectory are grouped with the ground truth placed on the left. In each group, a row shows the results corresponding to a loss function and a column results corresponding to a dataset.
4.5 Remark on computation cost
With a paritcle numbers around $2000$ to $4000$, the runtimes of an DPSI optimisation run that consists of $100$ iterations of forward-backward computations using $1$ datapoint can be done in about $10$ minutes. As the shown by Figures 4 and 5, most runs converge within $50$ iterations, which means that a plausible solutions can be found by DPSI in about $5$ minutes. With a good initial guess of the parameters, as shown by Figure LABEL: (to be updated), this runtime can be further reduced. In addition, the memory comsuptions for these simulations are within $3$ GB, which further permits implementations on portible and small GPUs. Therefore, this indicates that DPSI meets the computation costs criteria for practical deployment.
@@ 5 Conclusion
In addition, with the use of physics-based dynamics models, the identified parameters are physically meaningful. The experiments reveal that different parameter values found through optimisation can be interpreted in a way that is aligned with our understanding and intuitions about real-world physics. Therefore, the DPSI framework not only gives users confidence in the simulation controlled by the physics models and these parameters, but also provides users with intuitive angles to identify the limitations of the reconstructed manipulation dynamics.
@@ 5.1 Limitations and future research
Secondly, an important assumption of this work is that the deformation behaviour of the target material in need of parameter identification can be simulated by the selected elastoplasticity model. Thus, it should be noted that when the target material’s behaviour is largely underrepresented by the selected physics model, DPSI would not be able to produce realistic simulations. In other words, the manual selection of the elastoplasticity model is of high importance for the proposed DPSI framework. For example, the use of the von Mises plasticity model will make it difficult for DPSI to recover parameter values that accurately reflect realistic soil or foam behaviours, which are better captured by the Drucker–Prager model.
Thirdly, an important lesson learnt from the results is that increased complexity of motion and contact mode lead to less local minima and better system identification accuracy. The number of local minima indicates that the collected data is insufficient to fully induce the correct values of the concerned parameters. Thus, in the future, we envision a better framework that incorporates learning-based approaches to allow the automatic selection of diverse tools and interaction motions to achieve more efficient and accurate parameter identification.

### v2

@@ 4.4 Generalisation
(a) Triple-poking, object configuration 1
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(b) Triple-poking, object configuration 2
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(c) Flattening, object configuration 1
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(d) Flattening, object configuration 2
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(e) Poking-180-rotating, object configuration 1
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(f) Poking-180-rotating, object configuration 2
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 7: The long-horizon manipulation motions. Every subfigure visualises from top to bottom the real manipulation process and its simulations with the top two best DPSI-recovered parameters and hand-picked parameters.
@@
To recall, there are three unseen motions. The triple-poking motion moves the round end-effector to poke the object at three different locations along the $y$ axis. The flattening motion moves the cylinder end-effector to press and flatten the object towards the positive and negative $x$ directions. The poking-180-rotating motion moves the rectangle end-effector to press the object down and then rotate $180$ degrees about the $z$ axis. The triple-poking motion is less relevant to the friction coefficients, while the last two motions involve complex frictional contacts. Table 5 summarises the heightmap distances of the simulated results of the unseen motions (each with two object configurations). Figure 7 visualises the trajectories of the real manipulation trajectory and its simulations with the top two best DPSI-found parameters and the hand-picked parameters. The animation videos of them are included in Supplementary materials. Figure 8 presents the resultant particles and their heightmaps of each motion and object configuration simulated by the DPSI-found parameters. The following observations can be drawn.
@@
First, in Figure 7, (as well as the animation video), it can be seen that, for all motion and object configuration pairs, the real and simulated trajectories are highly similar. Figure 8 also shows that, although discrepancies exist, the simulated particles and their heightmaps of the manipulation results are also highly similar across the loss functions and datasets. This visual similarity proves that DPSI can recover physics parameters from simple interactions that allow the accurate simulation of unseen, longer, and more complex manipulation motions, exhibiting robustness against perception and data challenges. Moreover, we note that it is the single-datapoint cases that achieve $11$ out of $18$ top-three lowest heightmap losses across the three unseen motions as shown in Table 5, where $2$ achieved by the hand-picked parameters and $5$ by the 12mix and 6mix datasets.
Motion Object Num. datapoint 12 6 1 (rec.) 1 (round) 1 (cyl.) Avg. over losses
PCD CD 2384.72 2302.12 2394.64 3827.77 2288.37 2639.52
Config. PRT CD 3318.00 2845.03 2316.25 2309.38 4499.31 3057.59
1 PCD EMD 2700.57 2306.85 2971.97 3346.45 2255.47 2716.26
Poking-180 PRT EMD 2290.37 2299.37 3605.89 2361.34 2357.05 2582.81
Avg. over dataset 2673.42 2438.34 2822.19 2961.24 2850.05 -
Hand-picked parameters: *2132.04
rotating PCD CD 5367.10 5345.52 5473.18 4822.09 5770.72 5355.72
Config. PRT CD 4218.07 4227.91 5802.93 5766.13 *3994.08 4801.82
2 PCD EMD 4187.56 4469.51 4300.81 4026.08 5467.30 4490.25
PRT EMD 4175.07 4768.15 4323.36 4226.68 4978.39 4494.33
Avg. over dataset 4486.95 4702.77 4975.07 4710.25 5052.62 -
Hand-picked parameters: 5283.87
PCD CD 2956.06 2919.90 3028.85 3089.34 2949.11 2988.65
Config. PRT CD 2960.77 2971.96 2962.99 2963.06 *2798.67 2931.49
1 PCD EMD 2895.86 2875.01 2839.52 2863.56 3003.25 2895.44
Triple- PRT EMD 3100.87 3035.49 2852.95 3034.08 3059.61 3016.60
Avg. over dataset 2978.39 2950.59 2921.08 2987.51 2952.66 -
Hand-picked parameters: 2983.62
poking PCD CD *4175.07 4768.15 4323.36 4226.68 4978.39 4630.09
Config. PRT CD 4395.60 4399.77 4731.59 4685.70 4245.47 4491.62
2 PCD EMD 4430.33 4495.40 4417.59 4426.74 4379.40 4429.89
PRT EMD 4529.65 4304.59 4300.22 4272.90 4536.17 4388.70
Avg. over dataset 4508.76 4467.01 4499.59 4505.60 $4$.43 -
Hand-picked parameters: 4517.24
PCD CD 2443.54 2536.84 2153.18 2244.94 2403.77 2356.45
Config. PRT CD 2020.59 1988.33 2496.75 2547.35 1860.49 2182.70
1 PCD EMD 2149.35 *1835.63 2233.75 2196.98 2545.52 2192.25
Flattening PRT EMD 2275.80 2439.66 2407.77 2287.72 2539.93 2390.18
Avg. over dataset $2$.32 2200.12 2322.86 2319.25 2337.43 -
Hand-picked parameters: 2051.06
PCD CD 5325.90 5378.31 4752.87 4756.35 5743.95 5191.48
Config. PRT CD 5622.73 5387.05 5282.32 5119.63 5817.57 5445.86
2 PCD EMD 5058.42 4697.61 5798.36 5891.09 5190.65 5327.23
PRT EMD 4996.47 5460.32 5686.20 5152.61 5493.27 5357.77
Avg. over dataset 5250.88 5230.82 5379.94 5229.92 5561.36 -
Hand-picked parameters: *4483.75
(a) Object configure 1
(b) Object configure 2
Figure 8: The particles and heightmaps of the objects after applying the long-horizon unseen motions, simulated with the best set of physics parameters found at the level 2 contact complexity experiments. Darker colours of the heightmaps indicate greater heights of the object. For each motion, two object configurations are manipulated with three end-effectors. The results for each trajectory are grouped with the ground truth placed on the left. In each group, a row shows the results corresponding to a loss function and a column results corresponding to a dataset.
@@
Thirdly, the found solutions are not globally optimal. Despite the visual similarity, Table 5 shows that the best generalisation results are not always achieved by the cases with the best in-distribution validation performances, as the underlined and bolded texts rarely coincide. In addition, though the hand-picked parameters do achieve the best results in two out of six object simulations, they are not the best fit for every scenario. This again suggests the importance of re-fitting the parameters on the fly for every new scenario as there are hardly two objects that are identical especially for elastoplastic materials. Qualitatively, from the heightmaps and trajectories in Figures 8 and 7, one can also recognise certain visual discrepancies between the real objects and their simulation counterparts, indicating a certain degree of mismatch between the found parameters and the actual solutions. This could be improved by further diversifying the data (and the motions that produce them) used for optimisation. Other directions for improvement may include better observational noise removal or better physics model selection/approximation.
@@
Finally, the modelling limitations observed in the in-distribution visualisation still exist in the unseen motion visualisations (Figs. 7 and 8). The insufficient elastic returning at contact areas especially with sharp manipulator edges can be seen from all visualisations where the primary contact areas of the object rarely return to heights close to the real object. Floating displaced particles can be observed from the poking-180-rotating motion visualisations where many particles are carried away during rotations. The insufficient spreading of the influence of contacts can be recognised from the first object configuration of the flattening motion visualisations where the object is not tilted enough in the second half of the trajectory.
@@
In summary, the results and discussions above strongly support the argument that DPSI can recover physics parameters that produce the accurate simulation of unseen, longer, and more complex deformable object manipulation dynamics by optimising with very few real-world noisy and incomplete point-cloud data collected by simple and short interacting motions.
4.5 Cloud slime & soil materials
This subsection reports the DPSI performances with cloud slime and soil materials. Note that we use the same elastoplastic physics models in simulation. To do so, we collect a single interaction data point for each material using the cylinder end-effector and the poking-shifting-2 motion, optimise the parameters with the PRT EMD loss from a randomised initialisation for $100$ iterations with three random seeds, and validate the results on another data collected with the flattening motion and the cylinder end-effector. The real manipulation trajectory of the flattening motion and its simulations based on the found parameters are shown in Figure 9. The recovered parameters are presented in Table 6.
Figure 9 shows that DPSI manages to recover parameters that simulate cloud slime and soil as closely as possible despite the mismatch of the physics models and the target materials. Especially in the case of soil, Table 6 shows that DPSI finds large table friction coefficients and small end-effector friction coefficients to simulate the effect of the soil being flattened, although these are likely not close to the actual values. This suggests that DPSI will seek to find parameters that make the deformations of the simulated object as realistic as possible regardless of whether the physics models are a good match for the real materials. On the other hand, this also suggests that the manual selection of the physics models is important for DPSI to produce the best deformation simulation.
(a) Cloud slime
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
(b) Soil
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 9: Comparison of the real manipulation trajectories and the resultant height maps for the cloud slime and soil materials and their simulation counterparts based on the recovered physics parameters. In each subfigure, the second to last rows correspond to the results with three random seeds.
Random seed 1 2 3
Young’s modulus (kPa) 300000 106971 117831
Poisson’s ratio 0.4 0.33 0.03
Cloud Yield stress (kPa) 13836 3907 10469
slime Material density (kg/m3) 1000 1000 1491
Table friction 1.18 0.13 0.54
Manipulator friction 2.0 0.01 0.58
Random seed 1 2 3
Young’s modulus (kPa) 10000 88963 10696
Poisson’s ratio 0.32 0.33 0.08
Soil Yield stress (kPa) 9333 1000 10037
Material density (kg/m3) 2000 1653 1823
Table friction 2.0 1.08 1.83
Manipulator friction 0.01 0.01 0.01
Table 6: Physics parameters recovered by the DPSI method for Cloud slime and Soil materials with three random seeds.
4.6 Cause and treatment to local minima
As shown by the loss landscapes and the optimisation results, although DPSI can find good solutions, it could get stuck in local minima. Therefore, we performed extra experiments to optimise the PRT EMD loss with the three 1-datapoint datasets at contact level 1 with eight more random seeds to examine the random seed sensitivity of DPSI. The results are shown by the black dotted lines in Figure 10.
First of all, these dotted lines show consistent results with the previous conclusions: they converge to stable (though different) loss and parameter values, still reveal the negative relationship between the CD and EMD losses, and can achieve comparable performances with results from larger datasets. On the other hand, it can be seen that the main cause of converging to different solutions is the different initial parameter values, which were determined by the pseudo-random process controlled by different random seed values. This is not surprising as gradient-descent algorithms with complex physics problems are known to be sensitive to the choice of the initial solution (Hu et al., 2020; Antonova et al., 2023).
Figure 10: Validation losses (top five rows) and parameter values (last four rows) over $100$ gradient updates at level-1 contact complexity from optimising the loss functions with an extra set of $8$ random seeds. Solid lines (highly overlapped) are results with the same initial solution while dotted lines show results of optimising the PRT EMD loss with randomised initial solutions. Each column presents the results of optimising with a different dataset. Each row shows the changes of an evaluation metric or a parameter, denoted on the left. In each figure, different colours indicate the results of optimising a different loss, as labelled by the legend on the top.
To further test the assumption, we perform the same experiments with the four loss functions using the same set of seeds but manually initialise the solutions using the hand-picked parameters in Table 3 selected according to the material/geotechnical engineering literature (Rehman et al., 2018; Waheed and Asmael, 2023; StructX, 2014-2024). These hand-picked parameter values are less accurate in terms of validation loss as shown in Table 4. However, the (highly overlapped) solid lines in Figure 10 show that with this initial solution, DPSI converges with near-zero variances across eight random seeds, except for optimising the PCD CD loss with the 1-rec. dataset. This confirms that the differentiation process of the DPSI system is largely deterministic and the found solutions rely largely on the initial solutions, the optimisation algorithm and the loss functions. Moreover, with a good guess of the initial values, DPSI could converge very quickly (within $20$ iterations shown by the red/green solid lines).
Figure 11: Correspondence examples for the CD and EMD computations with hollow and filled point sets as targets. Colours indicate corresponding points. Note the CD loss is directional.
4.7 Remark on CD & EMD losses
As mentioned previously, the CD and EMD losses seem to focus on different spatial aspects of the object’s geometry, leading to deformations that match either the horizontal or the vertical directions. With the results of optimising the four losses with the same initial solutions shown in Figure 10, we can analyse the differences between CD and EMD losses with hollow or filled point sets.
Hollow or filled point sets: For the CD losses, it can be seen by comparing the blue and orange solid lines in Figure 10. The blue solids in the first two rows indicate that optimising the CD loss with a hollow point cloud target can achieve smaller losses in both CD loss metrics. However, the third to fifth rows in Figure 10 show that optimising CD with a filled particle system target (PRT CD loss) outperforms PCD CD in terms of both EMD loss metrics and the height map distance. On the other hand, optimising the EMD loss with a filled particle system target (red lines) exhibits the same pattern, consistently showing inferior performances in terms of the CD loss metrics while outperforming the PCD EMD losses (green lines) in terms of both EMD loss metrics and the height map distance. Therefore, the results indicate that optimising with a filled particle system target generated from a reconstructed watertight mesh is better than optimising with a hollow point cloud target.
The cause behind it is the inaccurate point-to-point correspondences generated with a hollow point set. As shown by Figure 11, a hollow point set either averages the spatial information of an area of particles over a thin layer of target points in computing the CD loss or ignores the information of a large body of particles within the particle body in computing the EMD loss. On the contrary, with a non-hollow target, the point correspondences connect spatially closer points and do not disregard information within the object body.
CD or EMD: Figure 10 shows that either with a hollow or non-hollow target, the EMD losses (green and red lines) consistently outperform the CD losses (blue and orange lines) in terms of the EMD and height map distance metrics while being inferior in terms of the CD metrics. Again, we believe that this is another consequence of the different ways CD and EMD compute the point correspondences.
Figure 11 shows that, with a hollow target, CD distributes the spatial information of the points within over the surface points, which is problematic. Take the shape in Figure 11 as an example, the particles above the target surface will be forced down, while the particles below the target surface will be pushed up, which is not supposed to happen. On the other hand, the EMD loss does not suffer from this effect because it corresponds the particles to the closest points from the target surface only. With a non-hollow target point set, the CD loss produces different correspondences in the two computation directions, which results in a mixture of correct and wrong spatial relations, while the EMD loss is more likely to find the best match spatially due to the one-to-one correspondence generated by a linear assignment algorithm.
In short, this analysis leads to the conclusion that the PRT EMD loss, which computes point correspondences between the filled simulation particle system and a non-hollow target point cloud, is the most ideal choice for capturing the correct spatial differences between point sets.
4.8 Remark on computation cost
With particle numbers around $2000$ to $4000$, a DPSI optimisation run that consists of $100$ iterations of forward-backward computations using $1$ datapoint can be done in about $10$ minutes. As shown by Figures 4 and 5, most runs converge within $50$ iterations, which means that a plausible solution can be found by DPSI in about $5$ minutes. With a good initial guess of the parameters, as shown by the dotted lines in Figure 10, this runtime can be further reduced to around $2$ minutes. In addition, the memory consumption for these simulations is within $3$ GB, which permits implementations on portable and small GPUs. With no more than $2$ minutes for the 3D reconstruction pipeline, a parameter identification process could take less than $5$ minutes to complete. These statistics confirm that DPSI meets the criteria in computational costs for practical deployment. It also suggests again that it is practical to perform parameter identification with DPSI at any time without the need to rely on previously found parameters or slow manual measuring techniques.
@@ 4.8 Remark on computation cost
Table 7: Computational costs (time and GPU memory consumption) of the TaiChi-based differentiable simulator used in the DPSI framework with a Nvidia GeForce RX 4090 GPU. The costs of simulations with three motions and particle densities are reported. CP.: compile-time costs. The first time a program is run takes extra time to compile TaiChi kernels. $100$ FB-iterations (min): the runtime of $100$ forward and backward computations in minutes.
@@ 5 Conclusion
In addition, with the use of physics-based dynamics models, the identified parameters are physically meaningful. The experiments reveal that different parameter values found through optimisation can be interpreted in a way that is aligned with our understanding and intuitions about real-world physics. Therefore, the DPSI framework not only gives users confidence in the simulation controlled by the physics models and these parameters but also provides users with intuitive angles to identify the limitations of the reconstructed manipulation dynamics.
@@ 5.1 Limitations and future research
Secondly, an important assumption of this work is that the deformation behaviour of the target material in need of parameter identification can be simulated by the selected elastoplasticity model. Thus, it should be noted that when the target material’s behaviour is largely underrepresented by the selected physics model, DPSI would not be able to produce very realistic simulations. In other words, the manual selection of the elastoplasticity model is of high importance for the proposed DPSI framework. For example, our results with soil reveal that DPSI found parameters to make simulations appear as close to the real materials as possible, but the use of the von Mises plasticity model will make it difficult to recover parameter values that accurately reflect realistic soil or foam behaviours, which are better captured by the Drucker–Prager model.
Thirdly, an important lesson learnt from the results is that increased complexity of motion and contact mode leads to fewer local minima and better system identification accuracy. The number of local minima indicates that the collected data is insufficient to fully induce the correct values of the concerned parameters. Thus, in the future, we envision a better framework that incorporates learning-based approaches to allow the automatic selection of diverse tools and interaction motions to achieve more efficient and accurate parameter identification.